OpenStudy (superhelp101):

Classify the conic section and write it's equation in standard form. -x^2 + 10x + y -21=0 This is a parabola but I don't know why and also I need help on making it into standard form.

OpenStudy (superhelp101):

@turningtest

OpenStudy (jhannybean):

It's a parabola because you've got an \(x^2\) term, which implies a quadratic equation.

OpenStudy (superhelp101):

@turingTest

OpenStudy (superhelp101):

My teachers got y=(x-5)^2 -4

OpenStudy (superhelp101):

I don't really understand the steps

OpenStudy (superhelp101):

Actually Ohhh u completed the square

OpenStudy (anonymous):

Put all the y's on one side, and all the x's and constants on other side

OpenStudy (anonymous):

Then do some factoring

OpenStudy (jhannybean):

Hmm,let's try this agani.

OpenStudy (jhannybean):

Oh, I see. \[-x^2 + 10x + y -21=0\implies -y = -x^2 +10x-21 \implies y=x^2 -10x+21\]

OpenStudy (jhannybean):

Then... \[y=x^2-10x+21\]\[y=(x^2-10x)+21\]\[y=(x^2-10x+25-25)+21-25\]\[y=(x-5)^2 -4\]

OpenStudy (jhannybean):

I completed the square to find my parabola in vertex form.

OpenStudy (superhelp101):

So did we basically complete the square twice in the problem?

OpenStudy (jhannybean):

Twice? The first time I messed up.

OpenStudy (jhannybean):

Just look at the last portion :)

OpenStudy (superhelp101):

Oh alright that makes sense! Thank you! If I have any other questions I will ask u :)

OpenStudy (jhannybean):

Sure thing :) I will try.